OpenStudy (anonymous):

how to calculate the molecular weight if the molar volume and the specific volume are given

OpenStudy (anonymous):

Consider specific volume is measured in units of:\[\frac{m^3}{kg}\] Molar Volume is measured in \[\frac{m^3}{mol}\] You want kg/mol so you want: \[\frac{molar weight}{specific weight}\] This gets you \[\frac{kg}{mole}\] You may need to play with units to get the desired units i.e. g/mol.

OpenStudy (anonymous):

It you write out the units for molar weight and divide them by the units for specific weight you will see the m^3 cancel. Then times the units by kg/kg and mol/mol to get the units in the correct orientation.

OpenStudy (anonymous):

I get it thanks

OpenStudy (anonymous):

gotta a problem with the units

OpenStudy (anonymous):

a sample has a specific volume of 5.2 (m^3)(1/kg) and a molar volume of 1160 (m^3)(kg*1/mol) determine the molecular weight of the material

OpenStudy (anonymous):

I think they may have supplied incorrect units, see: http://en.wikipedia.org/wiki/Molar_volume. \[\frac{1160\frac{m^3}{mol}}{5.2\frac{m^3}{kg}}\frac{kg}{kg}\frac{mol}{mol}=2.2*10^2\frac{kg}{mol}\]